OpenStudy (anonymous):

What is the slope of the line whose equation is –48 = 2x − 8y?

OpenStudy (alexandervonhumboldt2):

first convert into sloe intercept form: first subtract 2x from both sides

OpenStudy (anonymous):

8y ?

OpenStudy (alexandervonhumboldt2):

-48-2x=2x-8y-2x -8y=-48-2x now divde both sides by -8: -8y/-8=(-48-2x)/-8 y=(48+2x)/8=?